Solve this system of equations by the substitution method.

x+4x=-1
3x+y=8

Found 2 solutions by Jstrasner, MathTherapy:
Answer by Jstrasner(112)

Hey,
So for this one, we simply solve the first one:
x+4x=-1 => 5x=-1 => x=-1/5
then we substitute this into the next equation:
3(-1/5) + y = 8
And solve for this one:
-3/5 + y = 8 => y= 43/5
I hope this helps!

Answer by MathTherapy(10858)

Solve this system of equations by the substitution method.
 
x+4x=-1
3x+y=8
***************
The other person's solutions for x and y are correct, considering that the 1st equation of the system is as stated. 

However, it is this author's strong belief that it's missing the variable y, and so, it's quite foolish, in this author's opinion,
to guess what the true and correct equation is, and come up with the correct answer. That's asking too, too much, for sure!!

It's on the person asking for help to clearly state the problem, a fact that is most likely stated somewhere on this site.

How else will he/she expect to get a correct answer?